A rectangle scale used to weigh an elephant seal is 8 feet longer than it is wide. The area of the scale is 33 square feet. What are the dimensions of the scale?
Answer by Tutor Online(2)

A rectangle scale used to weigh an elephant seal is 8 feet longer than it is wide. The area of the scale is 33 square feet. What are the dimensions of the scale?
Let width be x feet
Then length would be x+8 , as its given scale is 8 feet longer then its width
Area of rectangle is given by
Area= length * width
Area of scale = x(x+8) sq feet
x ( x+8 ) = 33
x^2 +8x -33 =0
Solving above quadratic by factorization
x^2 + 11x -3x -33 =0
x(x+11) -3 ( x+11)=0
(x+11)(x-3) =0
So either x = 3
x = -11
x cannot be negative as width cannot be negative so x = 3 is the answer
So width of scale is 3
hope this helps